OpenStudy (anonymous):

The radius of Venus is 3,760.4 miles. What is the approximate volume of Venus?

OpenStudy (arabpride):

Oh the irony :'D Wutchu tryin to say-- i'm PHAT?!?!?!??!?!!! x'DDD

OpenStudy (arabpride):

But yah okay so the volume of a sphere = 4/3*pi*r^3

sammixboo (sammixboo):

\(\rm \Large \color{red}{~Volume~Of~A~Sphere}\) V = \(\rm \frac{4}{3}\) * 3.14 * \(\rm r^3\) V = \(\rm \frac{4}{3}\) * 3.14 * \(\rm 3,760.4^3\)

sammixboo (sammixboo):

3,760^3 doesn't work on the calculator! Ahaha

sammixboo (sammixboo):

Wait never mind

OpenStudy (arabpride):

You know - I always just use google right?! All you have to put in is "volume of...." into the search bar -and you'll get a calculator to help you solve for the volume of whatever it is~ cx

sammixboo (sammixboo):

First we do the exponent. \(\rm 3760.4^3\) 3760.4 * 3760.4 * 3760.4 = 53174342924.9 Now we plug in 53174342924.9 where \(\rm 3760.4^3\) was

sammixboo (sammixboo):

V = \(\rm \frac{4}{3}\) * 3.14 * 53174342924.9

sammixboo (sammixboo):

Whew this problem

sammixboo (sammixboo):

Now you multiply \(\rm frac{4}{3}\) by 3.14 and get 4.81666667 Now you plug in 4.81666667 where \(\rm \frac{4}{3}\) * 3.14 was

sammixboo (sammixboo):

4.81666667 * 53174342924.9

sammixboo (sammixboo):

SO what is 4.81666667 * 53174342924.9

sammixboo (sammixboo):

Ugh now he went offline Imma finish the problem anyways

sammixboo (sammixboo):

4.8166666 * 53174342924.9 = 256123081543

sammixboo (sammixboo):

That is your answer

sammixboo (sammixboo):

I hope

sammixboo (sammixboo):

LOL VAE your volume is 256123081543 something

OpenStudy (anonymous):

4/3 * (3760.4)^3

TheSmartOne (thesmartone):

Wow and no one told him Wrong subject and Welcome to Open Study!!

TheSmartOne (thesmartone):

You are huge @arabpride lol

OpenStudy (arabpride):

x'D I know i am~